OpenStudy (anonymous):

X^2 - 5x - 36 + F(x)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

f(x) =x^2 - 5x - 36 [Which factors of -36 sum to -5?] -9 x 4 Substitute for -5x f(x) =x^2 - 9x + 4x- 36 [factorise in parts] f(x) =(x^2 - 9x) + (4x- 36 ) f(x) =x(x - 9) + 4(x- 9 ) f(x) =(x+ 4) (x - 9) Hence x = -4 or x = 9